# OSS Malicious Package Analysis in the Wild

### v2

@@ IV-E Attack Campaign Evolution (RQ4)
Operations Evolution. We provide 5 kinds of operations: (1) changing name (CN); (2) changing version (CV); (3) changing description (CD); (4) changing dependency (CDep); and (5) changing source code (CC). A changing operation is mapping to any of the five, {CN, CV, CD, CDep, CC}. Figue 12 plots the distribution of those operations among the malware groups, where the X-axis is the operation type, and the Y-axis is the percentage. We observe that CN is the most popular operation, and CV and CDep are the least popular operations. The CN’s percentage is 98.92% rather than 100%. This indicates that 1.088% (100%- 98.92%) release attempts do not use a different package name because old malicious packages are not detected by the registry manager. In this case, attackers would use a different version to make malicious packages look more like legitimate ones. The CC’s percentage is 39.76%, indicating that source code changing is a common operation. We manually inspected the content of the CC operations and found that the line of changing code was around 3.7 lines. This indicates malicious packages with CC operations still use similar attack behaviors.
IDN | Operation | IDN | Operation |
66,092,932 | (CDep, CD, CN, CC) | 83,945 | (CD, CN, CC) |
240,819 | (CDep, CD, CN, CC) | 28,309 | (CD, CN) |
142,040 | (CD, CN, CC) | 21,635 | (CDep, CD, CN, CC) |
89,379 | (CDep, CD, CN, CC) | 20,431 | (CDep, CD, CN, CC) |
89,131 | (CN) | 17,888 | (CDep, CN, CC) |
Further, we combine the download number and operation to reveal the malware’s impact. Table VIII lists the top 10 of increasing download numbers (IDN) and corresponding operations. The IDN is equal to the difference between and . A large IDN reflects the great impact of a changing operation. We observe that the top 10 IDNs are following the process: (1) first releasing an initial package without any malicious behaviors; (2) constantly adding new features, and (3) continuing to camouflage it as a popular package until detected by the registry manager. From the attacker’s perspective, the multi-faceted changing operation may expand the malicious package’s impact range.
Insights that we have learned are as follows. (1) Although the registry manager quickly removes malicious packages, the attacker releases the malware repeatedly. The impact of OSS malware is limited by a small download number. (2) The changing operation affects the download number and impacts.
V Related Work
@@
Malware Analysis is to obtain a deep understanding of malware behaviors. Prior work on binary malware [11, 12, 13] leveraged the malware samples with traces to picture the malware behaviors. VX underground [14] is the largest open database for binary malware, which contains a large collection of malware samples and corresponding code bases. Rokon et al. [15] proposed SourceFinder to search the GitHub repository based on malware samples and provided a dataset of malware codes. Those databases are designed for binary malware, and they are not suitable for OSS malicious packages.
@@
In contrast, the research community has turned to OSS malware analysis, including the malware detection [66, 6, 7, 8, 9, 10], the malware behaviors [67, 68, 69], and the malware dataset [19, 70, 17]. Pfretzschner and ben Othmane [71] proposed a detection algorithm for dependency-based attacks on Node.js, and Staicu et al. [72] proposed a deep understanding of the injection attack on Node.js. Cao and Dolan-Gavitt [73] leveraged the artifacts to find 26 fork repositories among 68 popular cryptocurrency repositories. Several OSS malware datasets are presented in [16, 17, 70], whose package number is much smaller than ours.
@@
OSS Ecosystem is a complex network of software packages and developers. Decan et al. [74] and Zimmermann et al.[75] study that the package dependency has become more complicated as time passed, and the package security risk is also increasing, which can serve as a launching pad for attackers. Dann et al. [76] identified dozens or hundreds of the direct and transitive dependencies. Vasilakis et al. [77] proposed a vetting framework for reused components in internal/public repositories. Typosquatting is the most popular attack vector in the OSS ecosystem [78, 79]. Xiao et al.[80] proposed an attack that abuses hidden attributes, which attackers can exploit to obtain confidential data, bypass security checks, and launch denial-of-service attacks. Pashchenko et al.[81] claimed that many developers are unaware of dependency issues or do not want to modify the software. Zahan et al.[82] analyzed the metadata of 1.63 million packages and provided six indicators of security risks in the NPM ecosystem. Gu et al. [83] built RScouter and found 12 potential attack vectors in the 6 registries, which can be used to distribute malicious packages. Wyss et al. [84] proposed three SSC attacks, including install-time, import-time, and run-time attacks.
VI Conclusion
@@
In this paper, we build the largest OSS malware dataset from scattered online sources for curating a malware corpus in the research community. The dataset contains 23,000+ malicious packages, covering 200+ malware families and 10 OSS ecosystems, where security professionals can use it to develop and evaluate malware detection and analysis approaches. Further, we propose a knowledge graph to represent the OSS malware corpus and provide an empirical analysis of malicious packages in the wild. Our analysis reveals the diversity of malicious packages and the context of attack campaigns in the OSS ecosystem.

### v3

@@ V-B Malware Diversity (RQ2)
Social Engineering Tactics. Each subgraph is a group of multiple malicious packages with a similar code base. When a similar malware code is reused, its life cycle can be represented as {release detection removalchanging release }. The removal stage is not the end of the attack; instead, the same malware is released again in the subsequent phase, creating a loop. In each iteration, the malware must employ a social engineering tactic, as each OSS ecosystem has a security policy for taking down malicious packages. We use an operation to represent a social engineering tactic with multiple attack attempts. We use to represent a changing operation, where is the difference between two packages, where is the prior malicious package and is the later package. We provide 5 kinds of operations: (1) changing name (CN); (2) changing version (CV); (3) changing description (CD); (4) changing dependency (CDep); and (5) changing source code (CC). A changing operation is mapping to any of the five, {CN, CV, CD, CDep, CC}. Figure 9 plots the distribution of those operations among the malware groups, where the X-axis is the operation type, and the Y-axis is the percentage. We observe that CN is the most popular operation, and CD and CDep are less frequent operations. The CN’s percentage is 88.65%, and the CV’s percentage is 11.35%. Hence, the CV operation indicates that prior malicious packages are not taken down by the registry. In this case, attackers would use a different version to make malicious packages look more like legitimate ones. Once a malicious package is taken down from the registry, the same package name cannot be reused. The CC’s percentage is 59.34%, indicating that the malware distributor will change the source code via a slight modification, e.g., use a different IP address. We manually inspected the content of the CC operations and found that the average number of lines changed was around 0.88 lines. The advantage of this social engineering tactic is that it is simple and has low costs. The disadvantage is that today’s security tools easily detect this attack campaign. In short, the social engineering tactic for OSS malware is simple: using a different name; using a different version when the malware is not detected; and a slight change in code but keeping the original attack behavior and functions.
Active Time. Each subgraph has a starting time to distribute its first malware and the last time to distribute its last malware. We use to denote the timestamp of the first malware and to denote the timestamp of the last malware. The time span between and represents the active period of the repeated attack, denoted as . The active period is used to represent the time span of the subgraph, where the malware is active intermittently in the OSS ecosystem. Figure 10 plots the CDF of the active period of the repeated attack. It is observed that 80% of the subgraph have an active time of less than 15 days. This shows that most repeated attacks are achieved by sending a large number of similar packages within a short time frame. In addition, there are still 53 subgraphs with time intervals of more than 60 days, some of which span over 1,000 days, indicating that a similar code base has been repeatedly used over this period.
V-C Dependent-hidden (RQ3)
The key point of dependent-hidden is that attackers use a different package to camouflage or hide malicious packages to evade security detection tools. We use dependency subgraphs to represent OSS malware: if two nodes have a dependency edge , we put them into the same subgraph. Table VII lists the overall dependent-hidden subgraph, including 22 groups in NPM, 13 groups in PyPI, and 3 groups in RubyGems. We find one large subgraph in 3 OSS ecosystems (the maximum size in NPM is 232, the maximum size in PyPI is 950, and the maximum size in RubyGems is 34). The remaining subgraphs have a small size (less than 10). The largest subgraph is formed by multiple dependencies of packages that are reused by different malicious packages.
OSS | Pkg. Num.# | Subgraph Num.# | Ave. Num.# |
NPM | 323 | 22 | 14.68 |
PyPI | 992 | 13 | 76.31 |
RubyGems | 39 | 3 | 13 |
OSS | Dependency (Num. of other malicious packages) |
NPM | util(88), icons(39), common(4), object-color(3), settings(3) |
PyPI | urllib(448), request(124), urllib3(92), timedelta(75) |
values(18), public(14), pystyle(12), urlsplit(12), coloram(11) | |
pwd(11), connection(10), pkgutil(10), twyne(8), runcmd(8) | |
docutils(6), seccache(6), openvc(5), faq(4), setupcfg(4), | |
exit(4), load(3), jsfiddle(3) | |
Ruby | rest-client(32) |
We further inspect the details of dependent-hidden subgraphs. Table VIII shows the malicious packages that were reused by more than two different malicious packages. In the NPM ecosystem, the most frequently referenced package is ‘util’, which was claimed as a dependency library by 88 different malicious packages. In the PyPI ecosystem, ‘urllib’ was referenced 448 times, while in the RubyGems ecosystem, ‘rest-client’ was used as a dependency of 32 different malicious packages. Combining Table VII and VIII, we observe that 1,354 malicious packages used 28 malicious packages as their dependency library. In addition, the dependent-hidden subgraphs have two methods to hide malicious packages, as follows. (1) Attackers tend to use common names often found in software development, such as ‘util(s)’, ‘common’, and ‘setting’, as malicious package names. (2) Some popular legitimate packages may be embedded with malicious code in a specific version.
Active Time. Figure 11 plots the CDF of the active period of dependent-hidden subgraphs. 80% dependent-hidden groups have less than 10 days of active periods. It is caused by a different attack strategy. A dependent-hidden subgraph indicates that the attacker releases a package with malicious code and then creates a second one, depending on the first one. Compared with similar subgraphs (45.16 days on average), the dependent-hidden subgraph has a shorter average active period (10.5 days). The reason is that a quick and simple attack strategy may be effective in deceiving users into downloading and installing malware. We also find a long-tail effect of the active period for those attack campaigns, indicating some malicious packages have been released after more than 100 days. In short, the costs of attacking OSS ecosystems by exploiting dependencies are high, while the benefits are low.
OSS | Pkg. Num.# | Subgraph Num.# | Ave. Num.# |
NPM | 3,110 | 33 | 94.24 |
PyPI | 7,249 | 40 | 181.23 |
RubyGems | 76 | 9 | 8.44 |
V-D Malware Context (RQ4)
As we mentioned before, security reports disclose the information about corresponding SSC attack campaigns. We use the co-existing subgraphs to represent OSS malware: if two nodes have a co-existing edge , we put them into the same subgraph. Table IX lists the overall co-existing subgraphs, including 33 subgraphs in NPM, 40 subgraphs in PyPI, and 9 subgraphs in RubyGems. The co-existing subgraph in PyPI contains 181.23 malicious packages on average; the average size in NPM is 94.24; and the average size in RubyGems is 8.44. A security report may contain multiple malicious packages, indicating a many-to-one mapping between a report and packages.
Social Engineering Tactics. Figure 12 depicts the operation distribution of subgraphs. We observe that CN is the most popular operation, and other operations have minimal size. The CN’s percentage is 94.83%. Subsequently, the CV’s percentage is 5.17%, and the CC’s percentage is 5.28%. It is evident that Figures 9 and 12 are similar in the distribution of operations. The reason is that many malicious packages from security reports belong to the malware based on similar code.
Active Time. Figure 13 plots the CDF of the active period of subgraphs in security reports. We observe that 20% of the reported attacks have an active time of 0 days. It is the reason that those malicious packages were released and taken down at the same time. We find that the curve rises slowly, indicating that attack campaigns cover various active times. This is because commercial corporations and security professionals publish the newest SSC attack events as soon as possible. For example, Tianwen [22] releases a security per quarter.
Context. The rationale behind the reported attack is that security reports contain context about OSS malicious packages. Security analysis reports disclose many types of attack activities. Analysis reports can supply relevant contextual information about attack activities, e.g., the attacker information and attack method.
We extract the relevant context from the report from the subgraphs. Specifically, we use the Indicator of Compromise (IoC) as the attack context, which is the warning sign that a system may have been breached or compromised. Here, there are 3 types of IoC in the security reports, including suspicious IP addresses, malicious domains/URLs, and PowerShell commands/scripts. We counted the details of IoC indicators from our reported attack campaigns. There are 1,449 malicious URLs, 234 suspicious IP addresses, and 4 PowerShell commands. Note that those IoCs belong to the forensic data that security professionals use to detect potential malicious activity on a network or system.
We preprocess these IoCs, including removing the duplicate ones and identifying available ones. We explored IP addresses among 234. The maximum number of the same IP appears is 23 (46.226.*.*, 51.178.*.*, 81.24.*.*, 141.95.*.*, 135.181.*.*, 195.201.*.*, 5.135.*.*). We manually identify whether those IP addresses are available, and the results are that all IP addresses cannot be accessible. Figure 14 shows the top 10 domains among 1,449 URLs. The domain ‘bananasquad.ru’ has the largest number, 453 times, followed by ‘kekwltd.ru’ with 302 times. Those domains may belong to Command and Control Infrastructure, as C2, which is the server that attackers use to communicate with malicious packages for following initial exploitation. Malware uses them for remote control, data theft, or updating malicious instructions. We find that some URLs are used by malware to download executable files (such as .exe or .dll), or a malicious payload such as a script file or tool.
VI Security Implication and Application
This section demonstrates the security implications and applications of MalGraph.
VI-A Malware Detection
Algorithm | Acc. (w/o) | Acc. (w) | Recall (w/o) | Recall (w) |
RF | 0.897 | 0.944+5.2% | 0.825 | 0.984+19.3% |
LR | 0.841 | 0.859+2.1% | 0.806 | 0.836+3.7% |
KNN | 0.773 | 0.807+4.4% | 0.778 | 0.818+5.1% |
MLP | 0.860 | 0.895+4.1% | 0.839 | 0.927+10.5% |
As we mentioned before, many malicious packages are centered on a few aggressive packages that dominate the dataset. To build a classification model based on our dataset, we apply both machine learning and deep learning algorithms. In one category, we use the malware diversity to train the model; but in the other category, we do not use the malware diversity to train the model. Specifically, we evaluate the performance of three machine learning algorithms (Random Forest (RF) and Logistic Regression (LR), and K-Nearest Neighbors (KNN)) alongside one simple deep neural network (Multi-Layer Perceptron (MLP)).
The dataset includes all 3,840 tracked malware packages in NPM and 3,500 random legitimate packages from the registry. In each iteration, we randomly select a fixed number of clusters and sample two packages from each cluster into the test set, allowing repeated sampling for small clusters. The remaining samples are used to construct training sets in two ways: (1) we sample two packages from each cluster; (2) we randomly sample the same number of malicious packages from the remaining pool without using any clustering information. In both settings, we sample an equal number of legitimate packages for balance. We repeat this process for 50 iterations. We implement all classifiers using scikit-learn [57] and Pytorch [65]. Table X lists the average performance over 50 iterations of malware detection. With the help of malware diversity, the model achieves improvements in overall performance, with an average recall increase of around 10%. In short, MalGraph can improve the performance of malware detection.
VI-B Malware Analysis
Pkg. Num. | Malicious Behavior | |
NPM | 827 | Spyware, Backdoor, Data exfiltration via TLS |
414 | C2 channel, Credential collecting DNS tunneling exfiltration | |
196 | Beaconing, Fingerprint Spoofing C2 channel | |
149 | Webhook Abuse, Surveillance Data Exfiltration | |
118 | Webhook Abuse, Fingerprinting Service Abuse | |
118 | Beaconing, UA Spoofing, C2 channel | |
118 | Identity Exfiltration, Data Exfiltration | |
110 | Data Exfiltration, PII collecting OAuth2 Abuse | |
PyPI | 829 | Chinese Character Obfuscation Network Behavior Masking |
409 | Discord Payload Delivery, PowerShell Legitimate Package Spoofing | |
270 | Dropbox Malware Fetch, License Abuse | |
180 | Obfuscation, Legitimate Package Spoofing | |
140 | PowerShell, Legitimate Package Spoofing | |
134 | Dropbox Malware Fetch, PowerShell Legitimate Package Spoofing |
MalGraph can help security professionals to obtain a deep understanding of malicious behaviors. We provide a quantitative analysis of the malicious behaviors of OSS malware. Specifically, we leverage the similar edge in MalGraph to obtain OSS malware groups. If the group size is larger than 100, we inspect malware code to characterize its behaviors. The analysis approach is as follows. (1) If the malware is reported by online sources (Table III), we use the security report content to represent its behaviors. (2) If the malware is not reported, we use GPT-4o [66] to describe its behaviors. Then, we further inspect its code to avoid LLM’s hallucinations and errors. (3) All malicious behavior labels are curated by individual experience.
Table XI lists the quantitative analysis of malicious behaviors across NMP and PyPI. The results show that malicious packages in NPM primarily exhibit two behaviors: data exfiltration and obfuscation. Similarly, malicious PyPI packages disguise themselves as legitimate tools by using spoofed MIT licenses and development classifiers, evading detection with multi-layered obfuscation like Base64-encoded commands and hidden PowerShell windows. For example, attackers use unique characters and their combinations to achieve the obfuscation [67].
VII Discussion & Limitation
Malware Diversity plays a vital role in conventional malware research [68], guaranteeing the validity of the research. Each conventional malware has a lot of samples in binary form, and each sample is different from others at the binary level, called binary polymorphism. However, each OSS malware is released as a software package in the OSS ecosystem (without binary-level samples), including its metadata, dependencies, versions, and source code. The diversity of OSS malware is different from that of conventional malware, and we cannot use the malware family to represent it. So far, the diversity of OSS malware is unclear. It will be a future work to provide a new definition of the OSS malware diversity.
Missing Malware Packages. As we mentioned before, many malware packages are missing from our dataset. Once the malware is removed, there is a low probability that we can find it at other online sources. The only way to make amends for those malware packages is to cooperate between the industry and academia. We will release our malware corpus to the public, and we hope the community can help us supplement the missing packages.
Analysis Results Instability. One concern is that our analysis results may change over time when new malware packages are released or missing packages are supplemented. Indeed, the increasing number of malware packages may affect the result of the analysis. In our dataset, the release timeline of the malicious packages covers from 2014 to 2024. Hence, our analysis results are stable with time.
Data Representativeness. One concern is whether our dataset can reflect the whole OSS malware landscape. Due to distinct distribution channels, currently there lack of such a representative dataset. So far, our dataset has several advantages in acting as a representative dataset. (1) Our dataset involves the largest number of OSS malware compared with other existing datasets [16, 17, 69]. (2) We gather malware packages from a variety of online sources, ensuring a broad representation of the malware landscape. (3) We have cross-referenced our dataset with several prior studies to validate its relevance and comprehensiveness. (4) As shown in Figure 7, our dataset spans multiple years, encompassing both emerging and established threats, which strengthens its representativeness over time.
VIII Related Work
@@
Malware Analysis is to obtain a deep understanding of malware behaviors. Prior work on binary malware [11, 12, 13] leveraged the malware samples with traces to picture the malware behaviors. Bayer et al.[12] and Jiang et al.[13] proposed clustering techniques to partition malware samples into different groups based on their behaviors, and they analyzed the behavioral profiles of malware samples. VX underground [14] is the largest open database for binary malware, which contains a large collection of malware samples and corresponding code bases. Rokon et al.[15] proposed SourceFinder to search the GitHub repository based on malware samples and provided a dataset of malware codes. Those databases are designed for binary malware, and they are not suitable for OSS malicious packages.
@@
In contrast, the research community has turned to OSS malware analysis, including the malware detection [70, 6, 7, 8, 9, 10], the malware behaviors [71, 72, 73], and the malware dataset [18, 69, 17]. Pfretzschner et al. [74] proposed a detection algorithm for dependency-based attacks on Node.js, and Staicu et al. [75] proposed a deep understanding of the injection attack on Node.js. Cao et al. [76] leveraged the artifacts to find 26 fork repositories among 68 popular cryptocurrency repositories. Several OSS malware datasets are presented in [16, 17, 69], whose package number is smaller than ours.
@@
OSS Ecosystem is a complex network of software packages and developers. Decan et al. [77] and Zimmermann et al. [78] study that the package dependency has become more complicated as time passed, and the package security risk is also increasing, which can serve as a launching pad for attackers. Dann et al. [79] identified dozens or hundreds of the direct and transitive dependencies. Vasilakis et al. [80] proposed a vetting framework for reused components in internal/public repositories. Typosquatting is the most popular attack vector in the OSS ecosystem [81, 82]. Xiao et al. [83] proposed an attack that abuses hidden attributes, which attackers can exploit to obtain confidential data, bypass security checks, and launch denial-of-service attacks. Pashchenko et al. [84] claimed that many developers are unaware of dependency issues or do not want to modify the software. Zahan et al. [85] analyzed the metadata of 1.63 million packages and provided six indicators of security risks in the NPM ecosystem. Li et al. [86] proposed the dependency graphs between packages for analyzing known vulnerability impacts or underlying risks among different dependencies. Gu et al. [87] built RScouter and found 12 potential attack vectors in the 6 registries, which can be used to distribute malicious packages. Wyss et al. [88] proposed three SSC attacks, including install-time, import-time, and run-time attacks.
IX Conclusion
@@
In this paper, we build the largest OSS malware dataset from scattered online sources, including 24,000+ malicious packages. Further, we propose a knowledge graph to represent the OSS malware corpus, which security professionals can use to develop and evaluate malware detection and analysis approaches. Our quantitative analysis provides four insights into the research community: (1) OSS malware analysis research still needs to collect and clean the dataset; (2) the OSS malware diversity is an unsolved problem, causing noises and errors; (3) only 28 malicious packages were repeatedly hidden via dependency libraries; (4) the context of malware-based attacks are essential artifacts for today’s security community.
Acknowledgement
We are grateful to our shepherd Hiroshi Yamada and anonymous reviewers for their insightful feedback. The work was partially supported by the National Natural Science Foundation of China (No. 62272029 and No. 61972024).